# Orthogonal-by-construction augmentation of physics-based input-output models

### v1

@@ Abstract
Model augmentation is a promising approach for integrating first-principles-based models with machine learning components. Augmentation can result in better model accuracy and faster convergence compared to black-box system identification methods, while maintaining interpretability of the models in terms of how the original dynamics are complemented by learning. A widely used augmentation structure in the literature is based on the parallel connection of the physics-based and learning components, for both of which the corresponding parameters are jointly optimized. However, due to overlap in representation of the system dynamics by such an additive structure, estimation often leads to physically unrealistic parameters, compromising model interpretability. To overcome this limitation, this paper introduces a novel orthogonal-by-construction model augmentation structure for input-output models, that guarantees recovery of the physically true parameters under appropriate identifiability conditions.
@@
Physics informed and grey box model identification, Nonlinear system identification, Machine and deep learning for system identification
@@ 1 Introduction
In recent years, the increasing complexity of engineering systems and the growing performance demands in control applications have intensified the need for accurate nonlinear models. Discrete-time (DT) input-output (IO) models are one of the most commonly applied structures in system identification, as they incorporate a broad spectrum of model classes, ranging from linear time-invariant (LTI) (Hespanha, 2018) and linear parameter-varying (LPV) (Tóth, 2010) models to various nonlinear structures (Schoukens and Ljung, 2019). First-principle (FP) models can be obtained in DT IO form based on known physical laws and engineering insight. While such physics-based models provide interpretable system descriptions, they often capture only the dominant dynamics, while additional high-complexity effects, such as frictional properties or aerodynamic forces, are typically neglected.
@@
As an alternative approach, various data-driven identification methods have been developed for modeling nonlinear systems with an IO model structure (Schoukens and Ljung, 2019). In particular, recent advances employing deep artificial neural networks (ANNs) have demonstrated superior modeling accuracy compared to conventional approaches (Ljung et al., 2020). However, the practical use of ANN-based black-box models in control-oriented applications, e.g., trajectory planning, remains limited due to their lack of physical interpretability (Ljung, 2010). Furthermore, ANN-based models typically exhibit poor extrapolation capabilities beyond the range of the training data, and substantial learning effort is often spent on rediscovering system behaviors that are already well understood from first-principles knowledge.
@@
To address these challenges, different strategies have been proposed in the literature, starting from (light) grey-box modeling (Bohlin, 2006) till physics-informed neural networks (PINNs) (Raissi et al., 2019) and physics-guided neural networks (PGNNs) (Daw et al., 2022). One of the most promising directions of these hybrid approaches is model augmentation (Schön et al., 2022). The augmentation approach aims at combining FP models, i.e., baseline models, with flexible learning components to achieve faster convergence and better model accuracy compared to black-box learning methods (Djeumou et al., 2022). Furthermore, model augmentation produces interpretable models with a clear understanding of how the learning component complements the baseline dynamics.
@@
In this paper, we investigate a widely used model augmentation structure in the literature, namely the additive formulation. This approach connects the physics-based and learning components in parallel, and the corresponding parameters of the two components are jointly optimized. However, simultaneously tuning the learning-based and physical parameters results in the two ”subcomponents” competing with each other (Bolderman et al., 2022). As a result, the ANN can learn relations that could be represented by the FP model, while the baseline parameters can be tuned to physically unrealistic values. This effect undermines physical interpretability of the model estimate and can even compromise the extrapolation capabilities of the resulting model.
@@
This challenge was first addressed in Bolderman et al. (2022) in the context of PGNN-based feedforward control by introducing an additional regularization term into the cost function, penalizing deviations of the baseline parameters from their nominal values. This addition to the cost function effectively limits deviations of the baseline parameters compared to their initial values, hence preventing them from reaching a physically unrealistic parameter domain. Despite its simplicity, the approach has provided good experimental results (Bolderman et al., 2024); moreover, the method can be straightforwardly extended for more complex model augmentation structures (Hoekstra et al., 2025). Another attractive approach is based on an orthogonal projection-based regularization, introduced in Kon et al. (2022), also for feedforward control applications. This approach promotes a specific orthogonality between the baseline and learning components via regularization, penalizing when the ANN learns the already known relations represented by the baseline model. This technique has been adapted and generalized for nonlinear system identification in Györök et al. (2025); however, since these approaches promote orthogonality via regularization, inherently, there exists a trade-off between model accuracy and the desired complementarity. Finding the appropriate trade-off parameter (i.e., regularization coefficient) may not be intuitive. Therefore, we propose a direct parametrization with guaranteed orthogonality between the baseline and learning components on a selected data set without requiring any trade-off parameter. We also show that under certain conditions for the dataset used to impose such an orthogonality property, the tuned baseline parameters converge to their physically true values.
@@
  • Proving the consistency of the proposed model estimator and deriving that there is zero covariance between the estimated parameters corresponding to the baseline and the learning components.
@@
The remainder of the paper is organized as follows: Sect. 2 introduces the considered model augmentation problem with the additive structure and model learning setup. Then, Sect. 3 discusses the problems caused by the non-unique parametrization of the standard additive structure and proposes an orthogonal-by-construction model parametrization that addresses these challenges. In Sect. 4, the theoretical analysis of the proposed parametrization is presented. We provide conditions under which the presented model augmentation approach with joint parameter estimation recovers the physically true parameters of the baseline model, followed by the consistency analysis of the method. Sect. 5 shows a numerical example, where we demonstrate the effectiveness of the proposed orthogonal-by-construction model augmentation method. Finally, the conclusions on the achieved results are drawn in Sect. 6.
@@ 2 Problem statement
where $k\in\mathbb{Z}$ is the discrete time index, $y_{k}\in\mathbb{R}^{n_{y}}$ is the measured output, $x_{k}\in\mathbb{R}^{n_{x}}$ contains the lagged IO instances with $x_{k}=\mathrm{vec}(y_{k-n_{a}}^{k-1},u_{k-n_{b}}^{k})\in\mathbb{R}^{n_{x}}$, $y_{k-n_{\mathrm{a}}}^{k-1}=\begin{bmatrix}y_{k-1}^{\top}&y_{k-2}^{\top}&\dots&y_{k-n_{\mathrm{a}}}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{\mathrm{a}}n_{y}}$ being the lagged output values, and $u_{k-n_{\mathrm{b}}}^{k}$ can be defined similarly for the lagged input values. Moreover, $f:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}}$ is a nonlinear function, and $e_{k}\in\mathbb{R}^{n_{y}}$ is represented by a white noise process with finite variance. The formulation of (1) represents a wide range of systems depending on $n_{\mathrm{a}},n_{\mathrm{b}}\in\mathbb{R}$, e.g., NARX-type systems, when $n_{\mathrm{a}}>1$, $n_{\mathrm{b}}>0$.
@@
where $\hat{y}_{k}\in\mathbb{R}^{n_{y}}$ is the model output, $\theta_{\mathrm{b}}\in\mathbb{R}^{n_{\theta_{\mathrm{b}}}}$ contains the physical parameters, $\phi:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}\times n_{\theta}}$ is the regressor matrix-function of the baseline model. Furthermore, we assume that based on first-principles modeling, an initial rough estimate for the baseline parameters is available, and is denoted as $\theta_{\mathrm{b}}^{0}$.
@@
where $f^{\mathrm{ANN}}$ here represents a fully connected, feedforward neural network with $\theta_{\mathrm{a}}\in\mathbb{R}^{n_{\theta_{\mathrm{a}}}}$ being the collection of its parameters. Alternatively, $f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ can be replaced by any function approximator without loss of generality. Instead of the additive formulation, many other model augmentation structures can be selected from the literature, e.g., see Retzler et al. (2024). However, the additive, i.e., parallel, formulation offers a transparent model structure with clear separation between the baseline and learning components (Hoekstra et al., 2025), hence can be an attractive approach for practical applications.
To achieve the best possible data-fit, while simultaneously acquiring as accurate baseline parameters as possible, an efficient approach is to co-estimate $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters, as proposed in Bolderman et al. (2022). Hence, using a data sequence $\mathcal{D}_{N}=\left\{(x_{i},y_{i})\right\}_{i=0}^{N-1}$ generated by (1), the parameters are estimated by minimizing the prediction error loss function expressed as
@@
where $Y=\begin{bmatrix}y_{0}^{\top}&y_{1}^{\top}&\dots&y_{N-1}^{\top}\end{bmatrix}^{\top}$ are the stacked measured output values. Moreover, $\hat{Y}$ is computed, as
@@
where $\hat{Y}\in\mathbb{R}^{Nn_{y}}$ is the vectorized form of the model responses, $\Phi\in\mathbb{R}^{Nn_{y}\times n_{\theta_{\mathrm{b}}}}$ contains the baseline regressor matrices corresponding to the training data, while $F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}$ contains the learning component terms.
@@
where $\Phi\in\mathbb{R}^{Nn_{y}\times n_{\theta}}$ with $Nn_{y}>n_{\theta}$.
@@ 3.2 Direct parametrization of orthogonal subcomponents
The specified data informativity condition in Sect. 2 implies that $\Phi$ is full rank; hence, the inverse $\left(\Phi^{\top}\Phi\right)^{-1}$ exists. Moreover, $\left(\Phi^{\top}\Phi\right)^{-1}\Phi$ corresponds to the Moore-Penrose pseudo inverse of $\Phi$. With the introduced parameter $\theta_{\mathrm{aux}}$, the prediction map (5) is modified as
@@
$\Phi^{\top}F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}-\underbrace{\Phi^{\top}\Phi\left(\Phi^{\top}\Phi\right)^{-1}}_{I}\Phi^{\top}F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}=0.$ (12)
@@
For the applied parametrization, model training now results in the estimated $\theta_{\mathrm{b}}$, $\theta_{\mathrm{a}}$ parameters and, moreover, a fixed $\theta_{\mathrm{aux}}$ value. This is due to the applied orthogonal projection depending on the data distribution of $\mathcal{D}_{N}$. Hence, after training, prediction on new test data can be computed as
$\hat{y}_{k}=\phi(x_{k})\hat{\theta}_{\mathrm{b}}+f^{\mathrm{ANN}}_{\hat{\theta}_{\mathrm{a}}}(x_{k})-\phi(x_{k})\hat{\theta}_{\mathrm{aux}},$ (13)
where now $\hat{\theta}_{\mathrm{b}}$ and $\hat{\theta}_{\mathrm{a}}$ denote the estimated baseline and learning component parameters, while $\hat{\theta}_{\mathrm{aux}}$ is treated as a fixed parameter of the model.
@@ 4.1 Recovery of the baseline parameters
To analyze theoretically the error of the estimated baseline parameters, first, we reformulate the data-generating system as
@@
where $\delta:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{y}}$ represents the unmodeled terms, while $\theta_{\mathrm{b}}^{\ast}\in\mathbb{R}^{n_{\theta}}$ denotes the physically true baseline parameters. Then, a similar vectorized form of the true dynamics on the training data can be provided as in (5):
@@
where $Y\in\mathbb{R}^{Nn_{y}}$ is the vectorized form of the system outputs, and $\Delta\in\mathbb{R}^{Nn_{y}}$, $E\in\mathbb{R}^{Nn_{y}}$ are the collection of the $\delta(x_{k})$ and $e_{k}$ terms, respectively.
Achieving orthogonal subcomponents is only realistic if the baseline model regressor matrix function is in fact orthogonal to the unmodeled terms on a task-specific operating domain $x_{k}\in\mathbb{X}\subseteq\mathbb{R}^{n_{x}}$. Thus,
@@
For certain basis functions in $\phi$ and $\delta$ (e.g., orthogonal polynomials), (16) implies that with $N\to\infty$ there always exists an appropriate selection of regressor points $x_{k}$ to satisfy Condition 4. For other scenarios, a more detailed experiment design is necessary. An alternative interpretation is that (16) is the identifiability criterion of model class (9), while (17) is a specific excitation condition for the considered identification problem.
@@
Later, in Sect. 4.3, we will show that under certain conditions, Assumption 5 trivially holds with $N\to\infty$, i.e., we will prove consistency of the estimator. Now, we show that the proposed orthogonal parametrization recovers the physically true parameters of the baseline model.
@@
With Assumptions 1 and 5, Condition 4 holding, the estimation error of the baseline parameters is zero, i.e., $\hat{\theta}_{\mathrm{b}}\to\theta_{\mathrm{b}}^{\ast}$.
@@
$\Phi\theta_{\mathrm{b}}^{\ast}+\Delta=\Phi\hat{\theta}_{\mathrm{b}}+\left[I-\Phi\left(\Phi^{\top}\Phi\right)^{-1}\Phi^{\top}\right]F_{\hat{\theta}_{a}}^{\mathrm{ANN}}.$ (19)
@@
Re-arranging terms and taking the $\ell_{2}$ norm of both sides leads to
@@
4.2 Comparison without orthogonal parametrization
Two important factors can be derived from Theorem 6. First, when Assumptions 1 and 5 hold but Condition 4 does not, the true baseline parameters can not be recovered, but the acquired $\hat{\theta}_{\mathrm{b}}$ value is unique and the estimation error is given by (21). Secondly, using the standard additive structure (3) can result in larger baseline parameter errors (in a worst-case sense) than the orthogonal parametrization, even when Condition 4 is not satisfied. To show this, we make the following reasonable assumption.

### v2

@@ Abstract
This paper proposes a novel orthogonal-by-construction parametrization for augmenting physics-based input-output models with a learning component in an additive sense. The parametrization allows to jointly optimize the parameters of the physics-based model and the learning component. Unlike the commonly applied additive (parallel) augmentation structure, the proposed formulation eliminates overlap in representation of the system dynamics, thereby preserving the uniqueness of the estimated physical parameters, ultimately leading to enhanced model interpretability. By theoretical analysis, we show that, under mild conditions, the method is statistically consistent and guarantees recovery of the true physical parameters. With further analysis regarding the asymptotic covariance matrix of the identified parameters, we also prove that the proposed structure provides a clear separation between the physics-based and learning components of the augmentation structure. The effectiveness of the proposed approach is demonstrated through simulation studies, showing accurate reproduction of the data-generating dynamics without sacrificing consistent estimation of the physical parameters.
@@
Nonlinear system identification, Physics-based learning, Model augmentation
@@ 1 Introduction
In recent years, the increasing complexity of engineering systems and the growing performance demands in control applications have intensified the need for accurate nonlinear models. Discrete-time (DT) input-output (IO) models are one of the most commonly applied structures in system identification as they incorporate a broad spectrum of model classes, ranging from linear time-invariant (LTI) (hespanha_linear_2018) and linear parameter-varying (LPV) (toth_modeling_2010) models to various nonlinear structures (schoukens_nonlinear_2019). First-principle (FP) models can be obtained in DT IO form based on known physical laws and engineering insight. While such physics-based models provide interpretable system descriptions, they often capture only the dominant dynamics, while additional high-complexity effects, such as friction properties or aerodynamic forces, are typically neglected.
@@
As an alternative approach, various data-driven identification methods have been developed for modeling nonlinear systems with an IO model structure (schoukens_nonlinear_2019). In particular, recent advances employing deep artificial neural networks (ANNs) have demonstrated superior modeling accuracy compared to conventional approaches (ljung_deep_2020). However, the practical use of ANN-based black-box models in control-oriented applications, e.g., trajectory planning, remains limited due to their lack of physical interpretability (ljung_perspectives_2010). Furthermore, ANN-based models typically exhibit poor extrapolation capabilities beyond the range of the training data, and substantial learning effort is often spent on rediscovering system behaviors that are already well understood from first-principles knowledge.
@@
To address these challenges, different strategies have been proposed in the literature, starting from (light) grey-box modeling (bohlin_practical_2006) till physics-informed neural networks (PINNs) (raissi_physics-informed_2019) and physics-guided neural networks (PGNNs) (daw_physics-guided_2022). One of the most promising directions of these hybrid approaches is model augmentation (schon_multi-objective_2022; hoekstra_learning-based_2025). The augmentation approach aims at combining FP models, i.e., baseline models, with flexible learning components to achieve faster convergence and better model accuracy compared to black-box learning methods (djeumou_neural_2022). Furthermore, model augmentation produces interpretable models with a clear understanding of how the learning component complements the baseline dynamics.
@@
In this paper, we investigate a widely used model augmentation structure in the literature, namely the additive formulation. This approach connects the physics-based and learning components in parallel and the corresponding parameters of the two components are jointly optimized. However, simultaneously tuning the learning-based and physical parameters results in the two ”subcomponents” competing with each other (bolderman_feedforward_2022). As a result, the ANN can learn relations that could be represented by the FP model, while the baseline parameters can be tuned to physically unrealistic values. This effect undermines physical interpretability of the model estimate and can even compromise the extrapolation capabilities of the resulting model.
@@
This challenge was first addressed in bolderman_feedforward_2022 in the context of PGNN-based feedforward control by introducing an additional regularization term into the cost function, penalizing deviations of the baseline parameters from their nominal values. This addition to the cost function effectively limits deviations of the baseline parameters compared to their initial values, hence preventing them from reaching a physically unrealistic parameter domain. Despite its simplicity, the approach has provided good experimental results (bolderman_physics-guided_2024); moreover, the method can be straightforwardly extended for more complex model augmentation structures (hoekstra_learning-based_2025). Another attractive approach is based on an orthogonal projection-based regularization, introduced in kon_physics-guided_2022, also for feedforward control applications. This approach promotes a specific orthogonality between the baseline and learning components via regularization, penalizing when the ANN learns the already known relations represented by the baseline model. This technique has been adapted and generalized for nonlinear system identification in gyorok_orthogonal_2025; however, since these approaches promote orthogonality via regularization, inherently, there exists a trade-off between model accuracy and the desired complementarity. Finding the appropriate trade-off parameter (i.e., regularization coefficient) may not be intuitive. To overcome this, we propose a direct parametrization with guaranteed orthogonality between the baseline and learning components on a selected data set without requiring any trade-off parameter. We also show that under certain conditions for the dataset used to impose such an orthogonality property, the tuned baseline parameters converge to their physically true values.
@@
  • Proving that the proposed model estimator is consistent in the statistical sense and showing that there is zero covariance between the estimated parameters of the baseline and the learning components.
@@
The remainder of the paper is organized as follows: Sect. 2 introduces the considered model augmentation problem with the additive structure and model learning setup. Then, Sect. 3 discusses the problems caused by the non-unique parametrization of the standard additive structure and proposes an orthogonal-by-construction model parametrization that addresses these challenges. In Sect. 4, the theoretical analysis of the proposed parametrization is presented. We provide conditions under which the presented model augmentation approach with joint parameter estimation recovers the physically true parameters of the baseline model, followed by the consistency analysis of the method. Sect. 5 shows two numerical examples, where we demonstrate the effectiveness of the proposed orthogonal-by-construction model augmentation method. Finally, the conclusions on the achieved results are drawn in Sect. 6.
@@ 2 Problem statement
where $k\in\mathbb{Z}$ is the discrete time index, $y_{k}\in\mathbb{R}^{n_{\mathrm{y}}}$ is the measured output, $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ contains the lagged IO instances with $x_{k}=\mathrm{vec}(y_{k-n_{\mathrm{a}}}^{k-1},u_{k-n_{\mathrm{b}}}^{k})\in\mathbb{R}^{n_{\mathrm{x}}}$, where $y_{k-n_{\mathrm{a}}}^{k-1}=\begin{bmatrix}y_{k-1}^{\top}&y_{k-2}^{\top}&\cdots&y_{k-n_{\mathrm{a}}}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{\mathrm{a}}n_{\mathrm{y}}}$ for $n_{\mathrm{a}}>0$ and $\emptyset$ otherwise, being the lagged output values, and $u_{k-n_{\mathrm{b}}}^{k}$ can be defined similarly for the lagged input values. Moreover, $f:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ is a nonlinear function and $e_{k}\in\mathbb{R}^{n_{\mathrm{y}}}$ is represented by a white noise process with finite variance. The formulation of (1) represents a wide range of systems depending on $n_{\mathrm{a}}\geq 0$, $n_{\mathrm{b}}\geq 0$, called NARX-type systems.
@@
where $\hat{y}_{k}\in\mathbb{R}^{n_{\mathrm{y}}}$ is the model output, $\theta_{\mathrm{b}}\in\mathbb{R}^{n_{\theta_{\mathrm{b}}}}$ contains the physical parameters, $\phi:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{y}}\times n_{\theta}}$ is the regressor matrix-function of the baseline model. Furthermore, we assume that based on first-principles modeling, an initial rough estimate for the baseline parameters is available, and is denoted as $\theta_{\mathrm{b}}^{0}$.
@@
where $f^{\mathrm{ANN}}$ here represents a fully connected, feedforward neural network with $\theta_{\mathrm{a}}\in\mathbb{R}^{n_{\theta_{\mathrm{a}}}}$ being the collection of its parameters. Alternatively, $f^{\mathrm{ANN}}_{\theta_{\mathrm{a}}}$ can be replaced by any function approximator without loss of generality. Instead of the additive formulation, many other model augmentation structures can be selected from the literature, e.g., see retzler_learning-based_2024. However, the additive, i.e., parallel, formulation offers a transparent model structure with clear separation between the baseline and learning components (hoekstra_learning-based_2025), hence can be an attractive approach for practical applications.
To achieve the best possible data-fit, while simultaneously acquiring as accurate baseline parameters as possible, an efficient approach is to co-estimate $\theta_{\mathrm{b}}$ and $\theta_{\mathrm{a}}$ parameters, as proposed in bolderman_feedforward_2022. Hence, using a data sequence $\mathcal{D}_{N}=\left\{(x_{i},y_{i})\right\}_{i=0}^{N-1}$ generated by (1), the parameters are estimated by minimizing the prediction error loss function expressed as
@@
where $Y=\begin{bmatrix}y_{0}^{\top}&y_{1}^{\top}&\cdots&y_{N-1}^{\top}\end{bmatrix}^{\top}$ are the stacked measured output values. Moreover, $\hat{Y}$ is computed, as
@@
where $\hat{Y}\in\mathbb{R}^{Nn_{\mathrm{y}}}$ is the vectorized form of the model responses, $\Phi\in\mathbb{R}^{Nn_{\mathrm{y}}\times n_{\theta_{\mathrm{b}}}}$ contains the baseline regressor matrices corresponding to the training data, while $F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}$ contains the learning component terms.
@@
where $\Phi\in\mathbb{R}^{Nn_{\mathrm{y}}\times n_{\theta}}$ with $Nn_{\mathrm{y}}\geq n_{\theta}$.
@@ 3.2 Direct parametrization of orthogonal subcomponents
The specified data informativity condition in Sect. 2 implies that $\Phi$ is full rank; hence, the inverse $\left(\Phi^{\top}\Phi\right)^{-1}$ exists. In fact, $\left(\Phi^{\top}\Phi\right)^{-1}\Phi$ corresponds to the Moore-Penrose pseudo inverse of $\Phi$. With the introduced parameter $\theta_{\mathrm{aux}}$, the prediction map (5) is modified as
@@
$\Phi^{\top}F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}-\underbrace{\Phi^{\top}\Phi\left(\Phi^{\top}\Phi\right)^{-1}}_{I}\Phi^{\top}F_{\theta_{\mathrm{a}}}^{\mathrm{ANN}}=0.\vskip-12.0pt$ (12)
@@
For the applied parametrization, model training now results in the estimated $\hat{\theta}_{\mathrm{b}}$, $\hat{\theta}_{\mathrm{a}}$ parameters and, moreover, a fixed $\hat{\theta}_{\mathrm{aux}}$ value. This is due to the applied orthogonal projection depending on the data distribution of $\mathcal{D}_{N}$. Hence, after training, prediction on new test data can be computed as
$\hat{y}_{k}=\phi(x_{k})\hat{\theta}_{\mathrm{b}}+f^{\mathrm{ANN}}_{\hat{\theta}_{\mathrm{a}}}(x_{k})-\phi(x_{k})\hat{\theta}_{\mathrm{aux}}.$ (13)
@@ 4.1 Recovery of the baseline parameters
To analyze theoretically the error of the estimated baseline parameters, first, we reformulate the data-generating system (1) as
@@
where $\delta:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ represents the unmodeled terms, while $\theta_{\mathrm{b}}^{\ast}\in\mathbb{R}^{n_{\theta}}$ denotes the physically true baseline parameters. Then, a similar vectorized form of the true dynamics on the training data can be provided as in (5):
@@
where $Y\in\mathbb{R}^{Nn_{\mathrm{y}}}$ is the vectorized form of the system outputs, and $\Delta\in\mathbb{R}^{Nn_{\mathrm{y}}}$, $E\in\mathbb{R}^{Nn_{\mathrm{y}}}$ are the collection of the $\delta(x_{k})$ and $e_{k}$ terms, respectively.
Achieving orthogonal subcomponents is only realistic if the baseline model regressor matrix function is in fact orthogonal to the unmodeled terms on a task-specific operating domain $x_{k}\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$. Thus,
@@
For certain basis functions in $\phi$ and $\delta$ (e.g., orthogonal polynomials), (16) implies that with $N\to\infty$ there always exists an appropriate selection of regressor points $x_{k}$ to satisfy Condition 4. For other scenarios, a dedicated experiment design is necessary. An alternative interpretation is that (16) is the identifiability criterion of model class (9), while (17) is a specific excitation condition for the considered identification problem.
Remark 5
If the structure of $\delta$ can be inferred from prior physical knowledge, it is possible to specifically design an input sequence that satisfies Condition 4 (at least with respect to the suspected unmodeled terms). This is a common scenario in practical applications. For example, in aerial robotics, the drag force is known to depend quadratically on the relative velocity of the object (huang_leveraging_2024). However, the associated aerodynamic computations are often complex, and this force term is typically left unmodeled. Incorporating these physics-based insights can, however, guide the design of experiments that fulfil Condition 4.
@@
Note that Assumption 6 is only realistic for a finite $N$ number of data points if the training data does not contain any noise. Later, in Sect. 4.3, we will show that under certain conditions, Assumption 6 holds as $N\to\infty$ even when noise is present in the data, i.e., we will prove consistency of the estimator. Before moving to the noisy scenario, we show that the proposed orthogonal parametrization recovers the physically true parameters of the baseline model if the training data is noiseless.
@@
Under Assumptions 1 and 6, if Condition 4 is satisfied, then the estimation error of the baseline parameters is zero, i.e., $\hat{\theta}_{\mathrm{b}}=\theta_{\mathrm{b}}^{\ast}$.
@@
$\Phi\theta_{\mathrm{b}}^{\ast}+\Delta=\Phi\hat{\theta}_{\mathrm{b}}+\left[I-\Phi\left(\Phi^{\top}\Phi\right)^{-1}\Phi^{\top}\right]F_{\hat{\theta}\mathrm{a}}^{\mathrm{ANN}}.$ (19)
@@
Re-arranging the terms and taking the $\ell_{2}$ norm of both sides leads to
@@
4.2 Comparison with the standard parametrization
Two important factors can be derived from Theorem 7. First, when Assumptions 1 and 6 hold, but Condition 4 does not, the true baseline parameters can not be recovered, but the acquired $\hat{\theta}_{\mathrm{b}}$ value is unique and the estimation error is given by (21). Secondly, using the standard additive structure (3) can result in larger baseline parameter errors (in a worst-case sense) than the orthogonal parametrization, even when Condition 4 is not satisfied. To show this, we make the following reasonable assumption.